# MoMa: A Modular Deep Learning Framework for Material Property Prediction

### v2

@@ 2.3 Adaptive Module Composition & Fine-tuning
To better adapt to the downstream task $\mathcal{D}$, the composed module $g_{\mathcal{D}}$ is appended with a task-specific head and then fine-tuned on $\mathcal{D}$ to convergence.
@@ 3 Experiments
In this section, we conduct comprehensive experiments to demonstrate the empirical effectiveness of MoMa. The experimental setup is outlined in Section 3.1. The main results, discussed in Section 3.2, show that MoMa substantially outperforms baseline methods. Additionally, we conduct a thorough ablation study on the AMC algorithm as detailed in Section 3.3. Confronted with the data scarcity challenge common in real-world materials discovery settings, we evaluate MoMa’s few-shot learning ability in Section 3.4, where it achieves even larger performance gains compared to baselines. To further highlight the flexibility and scalability of MoMa, we extend MoMa Hub to include molecular datasets and present the continual learning results in Section 3.5. Finally, we visualize the module weights optimized by AMC in Section 3.6, highlighting MoMa’s potential for providing valuable insights into material properties.
@@ 3.1 Setup
To evaluate MoMa on material property prediction tasks, we conduct experiments on 17 tasks adhering to the benchmark settings established by Chang et al. (2022). Refer to Table 3 for more details.
@@
For the pre-trained backbone of MoMa, we choose to employ the open-source JMP model (Shoghi et al., 2023) for representing material systems given its superior performance in property prediction tasks across both crystals and molecules. For the evaluation metric, we report the average mean absolute errors (MAE) across five random data splits to enhance the robustness of the results. Additional implementation details, including the network architecture, the hyper-parameters for MoMa, and the computational cost, are provided in Section B.2.
@@
We compare the performance of MoMa with four baseline methods: CGCNN (Xie & Grossman, 2018), MoE-(18) (Chang et al., 2022), JMP-FT (Shoghi et al., 2023), and JMP-MT (Sanyal et al., 2018). CGCNN represents a classical method without pre-training. MoE-(18) trains separate CGCNN models for the upstream tasks of MoMa, then ensembles them as one model in a mixture-of-experts approach for downstream fine-tuning. JMP-FT directly fine-tunes the JMP pre-trained checkpoint on the downstream tasks. JMP-MT trains all tasks in MoMa with a multi-task pretraining scheme and then adapts to each downstream dataset with further fine-tuning. More discussions on baselines are included in Section B.3.
@@ 3.2 Main Results
As shown in Table 1, MoMa (Full) achieves the best performance with the lowest average rank of 1.35 and 14/17 best results. The adapter variant of MoMa follows, with an average rank of 2.47. Together, the two variants hold 16 out of 17 best results. They also exhibit the smallest rank deviations, indicating that MoMa consistently delivers reliable performance across tasks. Notably, MoMa (Full) outperforms JMP-FT in 14 tasks, with an impressive average improvement of 14.0%, highlighting the effectiveness of MoMa Hub modules in fostering material property prediction tasks. Moreover, MoMa (Full) surpasses JMP-MT in 16 out of 17 tasks with a substantial average margin of 24.8%, underscoring the advantage of MoMa in discovering synergistic knowledge modules.
@@
Among the baseline methods, JMP-FT performs the best with an average rank of 2.88, followed by JMP-MT with an average rank of 3.94. Though additionally trained on upstream tasks of MoMa Hub, JMP-MT still lags behind JMP-FT. We hypothesize that the inherent knowledge conflicts between the disparate material tasks pose a tremendous risk to the multi-task learning approach. We also observe that methods utilizing the JMP encoder outperform those based on CGCNN encoders. This demonstrates the good transferability of large force field models to material property prediction tasks.
3.3 Ablation Study of Adaptive Module Composition
Setup
We conduct a fine-grained ablation study of the Adaptive Module Fusion algorithm. The following ablated variants are tested: (1) Select average, which discards the weights optimized in Equation 3 and applies arithmetic averaging for the selected modules; (2) All average, which simple averages all modules in MoMa Hub; (3) Random selection, which picks a random set of modules in MoMa Hub with the same module number as AMC. Further analysis experiments are done using the MoMa’s full parametrization, i.e., MoMa (Full), due to its superior performance.
Figure 4: Ablation study of AMC. The main results using AMC  (purple) are compared with the ablated variants (orange) that substitute AMC with select average, all average and random selection. The axis represents the MAE on each dataset and smaller area is better. The ablated results are inferior to the main results in 13, 15 and 15 out of 17 tasks.
@@ Results
A visualization of the ablation results is presented in Figure 4. The ‘Select average’, ‘All average’, and ‘Random selection’ approaches perform worse to the main results using AMC in 13, 15, and 15 tasks, with an average increase of test MAE of 11.0%, 18.0%, and 20.2%. These results highlight the effectiveness of both the module selection and weighted composition strategies employed by AMC.
3.4 Performance in Few-shot Settings
@@
To better assess the performance of MoMa in real-world materials discovery scenarios, where candidates with labeled properties are costly to acquire and often exceptionally scarce (Abed et al., 2024), we construct a few-shot learning setting and compare the performance of MoMa with JMP-FT, the strongest baseline method. For each downstream task, we randomly down-sample $N$ data points from the train set to construct the few-shot train set, on which we apply the AMC algorithm to compose modules from MoMa Hub. Then we perform downstream adaptation by fine-tuning on the $N$ data points. The validation and test sets remain consistent with those in the standard settings to ensure a robust evaluation of model performance. Experiments are conducted with $N$ set to 100 and 10, representing few-shot and extremely few-shot scenarios.
@@ Results
The average test losses for the 17 downstream tasks of MoMa compared to JMP-FT across the full-data, 100-data, and 10-data settings are illustrated in Figure 5. As expected, the test loss increases as the data size decreases, and MoMa consistently outperforms JMP-FT in all settings. Notably, the performance advantage of MoMa is more pronounced in the few-shot settings, with the normalized loss margin widening from 0.03 in the full-data setting to 0.11 and 0.15 in the 100-data and 10-data setting. This suggests that MoMa may offer even greater performance gains in real-world scenarios, where property labels are often limited, thereby hindering the effective fine-tuning of large pre-trained models. Complete results are shown in Table 4.
Figure 5: The average test losses of MoMa and JMP-FT across 17 downstream tasks under varying data availability settings. MoMa consistently outperforms JMP-FT in all settings. The loss reduction amplifies as the data size shrinks, highlighting the advantage of MoMa in few-shot settings. Results are averaged over five random data splits.
3.5 Continual Learning Experiments
Motivation & Setup
Continual learning refers to the ability of an intelligent system to progressively improve by integrating new knowledge (Wang et al., 2024). We investigate this capability of MoMa by incorporating new modules into MoMa Hub. Due to its modular nature, it is expected that MoMa will exhibit enhanced performance in tasks that are closely aligned with the new modules, while maintaining its performance when these additions are less relevant. We expand MoMa Hub to include the QM9 dataset (Ramakrishnan et al., 2014) and evaluate the results across the 17 benchmark material property prediction tasks. For more details on the setup, please refer to Section B.4.
Results
We present the scatter plot of the reduction rate of test MAE w.r.t. the proxy error decrease in Figure 6 across datasets where QM9 modules are selected. We observe that: (1) The integration of QM9 modules leads to an average of 1.7% decrease in test set MAE; (2) a larger reduction in the AMC-optimized proxy error correlates with greater performance improvements post-fine-tuning (with a Pearson correlation of 0.69). We highlight the task of MP Phonons prediction, which marks a significant 11.8% decrease in test set MAE following the expansion of MoMa Hub.
Figure 6: Scatter plot showing the relationship between the test MAE decrease and the proxy error (defined in Equation 3) decrease after the addition of QM9 modules. The dashed line represents the average test MAE decrease. The solid line fits the results with linear regression.
@@ 3.6 Materials Insights Mining
We contend that the AMC weights derived in Equation 3 can offer interpretability for MoMa as well as provide valuable insights into material properties. To explore this, we interpret the weights as indicators for the relationships between MoMa Hub modules and downstream tasks. Following Chang et al. (2022), we present a log-normalized visualization of these weights in Figure 7.
@@ Results
We make several noteworthy observations:
  • The weights assigned by AMC effectively capture physically intuitive relationships between material properties. For instance, the tasks of experimental band gap (row 1) and experimental formation energy (row 2) assign the highest weights to the computational band gap (columns 2 and 14) and formation energy modules (columns 1, 12, and 15). Also, for the task of predicting electronic dielectric constants, MoMa assigns high weights to the band gap modules, which is reasonable given the inverse relationship between the dielectric constant and the square of the band gap (Ravichandran et al., 2016).
@@
  • Some less-intuitive relationships also emerge. For the task of experimental band gap prediction (row 1), the formation energy module from the Materials Project (column 1) is assigned the second-highest weight. In the prediction of dielectric constant (row 9), modules related to thermoelectric and thermal properties (columns 5 and 6) are non-trivially weighted. However, the first-principles relationship between these tasks is indirect. We hypothesize that in addition to task relevance, other factors such as data distribution and size may also influence the weight assignments for AMC. Further investigation into these results are left to future work.
Figure 7: Heat map illustrating the AMC weights on one data split. The x-axis represents the task names of the MoMa Hub modules, while the y-axis shows the 17 material property prediction tasks in Table 1. Darker colors indicate higher weights, signifying a stronger correlation between the MoMa module and the downstream tasks.
4 Related Work
4.1 Material Property Prediction with Deep Learning
Deep learning methods have been widely applied for predicting material properties (De Breuck et al., 2021). In a seminal work by Xie & Grossman (2018), CGCNN is designed to model crystalline materials with multi-edge graphs and leverage graph neural networks to learn crystal representations. Since then, a series of research (Choudhary & DeCost, 2021; Yan et al., 2022; Das et al., 2023; Lin et al., 2023; Yan et al., 2024; Taniai et al., 2024) have focused on improving neural network architectures to better model the inductive biases of crystals for property prediction tasks.
Another line of work develops pre-training strategies to facilitate material property prediction (Jha et al., 2019; Magar et al., 2022; Zhang et al., 2023a; Wang et al., 2024; Song et al., 2024). Recently, a series of pre-trained force field models (Merchant et al., 2023; Batatia et al., 2023; Yang et al., 2024b; Neumann et al., 2024; Barroso-Luque et al., 2024) achieve remarkable accuracy in the stability prediction task of inorganic solid-state materials and show initial results in generalizing to a broader range of material properties. We highlight the JMP model proposed by Shoghi et al. (2023), which is trained on force and energy prediction tasks across multiple domains (small molecules, catalysts, etc.) and performs impressively when fine-tuned to downstream tasks of both molecules and crystals.
Extending beyond the prevailing pre-train and fine-tune paradigm, MoMa devises effective strategies to centralize material knowledge into modules and adaptively compose the modules to achieve superior downstream performance.
4.2 Modular Deep Learning
Modular deep learning (Pfeiffer et al., 2023) represents a promising paradigm in deep learning, where parameterized modules are composed, selected, and aggregated during the network training process. Different from the vanilla pre-train and fine-tune approach, modular methods employ composable network architectures that enable more tailored adaptations to different tasks and domains. Notable examples of modular networks include mixture-of-experts (Jacobs et al., 1991; Shazeer et al., 2016), adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021). Recently, we have seen an increasing number of successful applications of modular deep learning across domains such as NLP and CV (Puigcerver et al., 2020; Pfeiffer et al., 2020; Huang et al., 2023; Zhang et al., 2023b; Tan et al., 2024; Pham et al., 2024), where its strengths in flexibility and minimizing negative interference have been demonstrated.
In the field of material property prediction, the idea of modular deep learning is still under-explored. A work most similar to MoMa is proposed by Chang et al. (2022). Their framework, termed MoE-(18), integrates 18 models trained on various source tasks with mixture of experts. MoMa distinguishes itself from MoE-(18) in two key aspects: (1) MoE-(18) loads all pre-trained models indiscriminately for each downstream task, whereas MoMa adaptively composes a subset of relevant modules to mitigate knowledge conflicts and encourage positive transfer. (2) MoE-(18) is designed to address the data scarcity issue and is limited to the mixture-of-experts approach, while MoMa introduces modularity to target the inherent challenges in materials science and is not restricted to any specific modular method. Hence, MoMa marks the first systematic effort to devise a modular deep learning framework for materials.
5 Conclusion & Outlook

### v3

@@ 3.2 Adaptive Module Composition & Fine-tuning
Finally, to better adapt to the downstream task $\mathcal{D}$, the composed module $g_{\mathcal{D}}$ is appended with a task-specific head and then fine-tuned on $\mathcal{D}$ to convergence.
@@ 4 Experiments
In this section, we conduct comprehensive experiments to demonstrate the empirical effectiveness of MoMa. The experimental setup is outlined in Section 4.1. The main results, discussed in Section 4.2, show that MoMa substantially outperforms baseline methods. Additionally, we extend MoMa to more architectures in Section 4.3 and conduct an in depth examination of AMC in Section 4.4. Confronted with the data scarcity challenge common in real-world materials discovery settings, we evaluate MoMa’s few-shot learning ability in Section 4.5, where it achieves even larger performance gains compared to baselines. To further highlight the flexibility and scalability of MoMa, we extend MoMa Hub to include molecular datasets and present a scaling analysis of MoMa Hub in Section 4.6. Finally, we visualize the module weights optimized by AMC in Section 4.7, highlighting MoMa’s potential for providing valuable insights into material properties.
Table 1: Main results for 17 material property prediction tasks. The best MAE for each task is highlighted in bold and the second best result is underlined. The result for each task are the average of five data splits, reported to three significant digits. For each method, the standard deviation of the test MAE across five random seeds is shown in parentheses. Additionally, the average rank and its standard deviation across the 17 datasets are provided to reflect the consistency of each method.
Datasets CGCNN MoE-(18) UMA JMP-MT JMP-FT MoMa (Adapter) MoMa (Full)
Experimental Band Gap (eV) 0.471 (0.008) 0.374 (0.008) 0.355 (0.037) 0.377 (0.005) 0.358 (0.014) 0.359 (0.009) 0.305 (0.006)
Formation Enthalpy (eV/atom) 0.193 (0.015) 0.0949 (0.0016) 0.192 (0.020) 0.134 (0.001) 0.168 (0.007) 0.158 (0.009) 0.0839 (0.0013)
2D Dielectric Constant 2.90 (0.12) 2.29 (0.01) 2.34 (0.47) 2.25 (0.06) 2.35 (0.07) 2.31 (0.04) 1.89 (0.03)
2D Formation Energy (eV/atom) 0.169 (0.006) 0.106 (0.005) 0.120 (0.03) 0.140 (0.004) 0.125 (0.006) 0.112 (0.002) 0.0495 (0.0015)
Exfoliation Energy (meV/atom) 59.7 (1.5) 52.5 (0.8) 44.4 (11.5) 42.3 (0.5) 35.4 (2.0) 35.4 (0.9) 36.3 (0.2)
2D Band Gap (eV) 0.686 (0.034) 0.532 (0.008) 0.494 (0.061) 0.546 (0.020) 0.582 (0.018) 0.552 (0.014) 0.375 (0.006)
3D Poly Electronic 32.5 (1.1) 27.7 (0.1) 32.7 (6.0) 23.9 (0.2) 23.3 (0.3) 23.3 (0.2) 23.0 (0.1)
3D Band Gap (eV) 0.492 (0.008) 0.361 (0.003) 0.268 (0.016) 0.423 (0.004) 0.249 (0.001) 0.245 (0.002) 0.200 (0.001)
Refractive Index 0.0866 (0.0014) 0.0785 (0.0004) 0.0582 (0.0094) 0.0636 (0.0006) 0.0555 (0.0027) 0.0533 (0.0023) 0.0523 (0.0010)
Elastic Anisotropy 3.65 (0.11) 3.01 (0.03) 3.79 (2.48) 2.53 (0.26) 2.42 (0.36) 2.57 (0.61) 2.86 (0.28)
Electronic Dielectric Constant 0.168 (0.002) 0.157 (0.015) 0.116 (0.038) 0.137 (0.002) 0.108 (0.002) 0.106 (0.002) 0.0885 (0.0048)
Dielectric Constant 0.258 (0.008) 0.236 (0.002) 0.183 (0.034) 0.224 (0.004) 0.171 (0.002) 0.168 (0.002) 0.158 (0.002)
Phonons Mode Peak (cm-1) 0.127 (0.004) 0.0996 (0.0083) 0.0811 (0.0087) 0.0859 (0.0006) 0.0596 (0.0065) 0.0568 (0.0009) 0.0484 (0.0026)
Poisson Ratio 0.0326 (0.0001) 0.0292 (0.0001) 0.0225 (0.0014) 0.0297 (0.0003) 0.0221 (0.0004) 0.0220 (0.0003) 0.0204 (0.0002)
Poly Electronic 2.97 (0.10) 2.61 (0.13) 2.33 (0.89) 2.42 (0.03) 2.11 (0.04) 2.13 (0.03) 2.09 (0.03)
Poly Total 6.54 (0.24) 5.51 (0.04) 5.61 (1.49) 5.52 (0.03) 4.89 (0.06) 4.89 (0.04) 4.86 (0.07)
Piezoelectric Modulus 0.232 (0.004) 0.208 (0.003) 0.208 (0.027) 0.199 (0.002) 0.174 (0.004) 0.173 (0.003) 0.174 (0.001)
Average Rank 6.88 (0.33) 4.71 (1.40) 4.53 (1.42) 4.53 (1.28) 3.12 (1.54) 2.59 (1.12) 1.35 (0.86)
@@ 4.1 Setup
To better align with real-world material property prediction settings where labels are usually scarce, we conduct experiments on 17 low-data material property prediction tasks from Matminer (Ward et al., 2018) adhering to Chang et al. (2022). This benchmark offers a comprehensive evaluation of model capability on a wide span of properties critical for material discovery. Refer to Table 5 for more dataset details.
@@
For the pre-trained backbone of MoMa, we employ the open-source JMP model (Shoghi et al., 2024) for representing material systems given its superior performance in property prediction tasks across both crystals and molecules. For a rigorous comparison, we present the MAE averaged across the five splits adopted from Chang et al. (2022). Each experiment is repeated with five random seeds, and the reported standard deviation is computed across the seed-level averages. Additional implementation details, including the details of module architecture, the hyper-parameters for MoMa, and the computational cost, are provided in Section C.2.
@@
We compare the performance of MoMa with five baseline methods: CGCNN (Xie and Grossman, 2018), MoE-(18) (Chang et al., 2022), UMA (Wood et al., 2025), JMP-FT, and JMP-MT (Shoghi et al., 2024). CGCNN represents a classical method without pre-training. MoE-(18) trains separate CGCNN models for the upstream tasks of MoMa, then ensembles them as one model in a mixture-of-experts approach for downstream fine-tuning. UMA is a general-purpose atomic foundation model which achieves state-of-the-art performance in canonical benchmarks (Riebesell et al., 2023). We fine-tune the UMA-Medium model on each downstream task. JMP-FT directly fine-tunes the JMP pre-trained checkpoint on the downstream tasks. JMP-MT trains all tasks in MoMa Hub with a multi-task pretraining scheme and then adapts to each downstream dataset with further fine-tuning. More discussions on baselines are included in Section C.3.
@@ 4.2 Main Results
As shown in Table 1, MoMa (Full) achieves the best performance with the lowest average rank of 1.35 and 14/17 best results. MoMa (Adapter) follows, with an average rank of 2.59. Together, the two variants hold 16/17 best results. They also exhibit the smallest rank deviations, indicating that MoMa consistently delivers reliable performance across tasks. Notably, MoMa (Full) outperforms JMP-FT in 14 tasks, with an impressive average improvement of 14.0%, highlighting the effectiveness of MoMa Hub modules in fostering material property prediction. Moreover, MoMa (Full) surpasses JMP-MT in 16 of 17 tasks with a substantial average margin of 24.8%, underscoring the advantage of MoMa’s modular design in mitigating task interference. Further analyses in this section are done with MoMa (Full) due to its superior performance.
@@
Among the baseline methods, JMP-FT performs the best with an average rank of 3.12, followed by JMP-MT and UMA with an average rank of 4.53. Though additionally trained on upstream tasks of MoMa Hub, JMP-MT still lags behind JMP-FT. We hypothesize the inherent knowledge conflicts between disparate material tasks pose a tremendous risk to the multi-task learning approach. For UMA, it is primarily pre-trained on force-field datasets as a DFT surrogate, so its inductive bias may transfer less well to non-PES downstream tasks as compared to JMP.
4.3 Results with more architectures
To verify whether MoMa offers consistent benefits in other model backbones beyond JMP, we conduct additional experiments on the GNS architecture (Sanchez-Gonzalez et al., 2020) used by the Orb-v2 model (Neumann et al., 2024), which is not equivariant and much less complex than the GemNet-based architecture (Gasteiger et al., 2022) in JMP. Specifically, we first train and construct an Orb-based MoMa Hub. Then we run AMC and downstream fine-tuning identically as in Section 3.2. The results (Orb-MoMa) are compared with directly fine-tuning the pre-trained Orb model (Orb-FT). The average test MAE are reported on 5 splits and 5 random seeds.
As shown in Fig. 6(a), MoMa outperforms in 13/17 tasks and achieves a 6.1% average boost to direct fine-tuning. This indicates that the effectiveness of MoMa is consistent across GemNet-based and GNS-based architectures.
4.4 Ablation & Analysis of Adaptive Module Composition
Figure 6: (a) Results with Orb-v2; (b) Ablation study of AMC; (c) Analysis experiments of AMC. The axis represents test set MAE and smaller area is better.
Ablation Study
We conduct a fine-grained ablation study of AMC with three variants : (1) Select Average, which retains the AMC-selected modules (nonzero weights) but averages them uniformly; (2) All Average, simply averages all modules in MoMa Hub, which is equivalent to applying the classical Model Soup strategy (Wortsman et al., 2022); (3) Random Selection, which picks a random set of modules in MoMa Hub with the same module number as AMC. A visualization of the ablation results is presented in Fig. 6(b). The three variants are inferior to AMC in 13, 15 and 15 out of 17 tasks, with an average test MAE increase of 11.0%, 18.0% and 20.2%, respectively. This highlight the effectiveness of both module selection and weighted composition in AMC. The average test MAE of 5 splits are reported on one random seed (hereafter). Furthermore, we show in Section D.2 that AMC is robust to kkNN configurations and solver tolerances, yielding highly stable weights and consistent post–fine-tuning MAE.
Analysis Experiments
To empirically validate the benefit of AMC’s representation-driven and training-free pipeline, we replace AMC with three alternatives: (1) LoRAHub Learning (Huang et al., 2024), a black-box optimization approach for module composition; (2) JMP-(18), where we train a routing network over the 18 JMP MoMa modules; and (3) Softmax Weighting, a non-optimized heuristic based on kkNN proxy. As shown in Fig. 6(c), AMC consistently outperforms all baselines, surpassing the three variants on 15, 17, and 12 tasks with average MAE reductions of 21.8%, 15.5%, and 13.7%, respectively. This shows the benefit of the AMC over search-based, router-based and performance-based alternatives. See more details and discussion in Section C.5.
Efficiency Analysis
We highlight that AMC is highly efficient: it requires only a single round of forward embedding generation, followed by lightweight kkNN prediction and convex optimization. For the largest dataset, AMC converges in under 30 seconds. This efficiency enables MoMa to scale to a larger number of modules in future applications. See Section D.3 for a detailed analysis.
4.5 Performance in Few-shot Settings
Table 2: Few-shot evaluation. The average normalized test MAEs of MoMa and JMP-FT under varying data settings. MoMa consistently outperforms JMP-FT in all settings.
10-shot 100-shot Full data
JMP-FT 0.7003 0.4076 0.2217
MoMa 0.5503 0.2990 0.1871
Motivation & Setup
To better assess the performance of MoMa in real-world scenarios, where labeled material candidates are costly and often scarce (Abed et al., 2024), we construct a few-shot learning setting and compare MoMa with JMP-FT. For each downstream task, we down-sample the training data and apply AMC to compose modules from MoMa Hub, followed by fine-tuning on the sampled subset. The validation and test sets remain consistent with those in the standard setting for robust evaluation. Experiments are conducted under 10-shot and 100-shot conditions, representing few-shot and extremely few-shot scenarios.
@@ Results
The average normalized test MAEs [^1] Computed by dividing the test MAE of each task by its standard deviation. for the 17 downstream tasks of MoMa compared to JMP-FT across the full-data, 100-data, and 10-data settings are presented in Table 10. As expected, the test loss increases as the data size decreases, while MoMa consistently outperforms JMP-FT in all settings. Notably, the performance advantage of MoMa is more pronounced in the few-shot settings, with the normalized loss margin widening from 0.03 in the full-data setting to 0.11 and 0.15 in the 100-data and 10-data setting. This suggests that MoMa may offer even greater performance gains in real-world scenarios, where property labels are often limited, thereby hindering the effective fine-tuning of large pre-trained models. Complete results are shown in Section D.4.
4.6 Scaling Analysis of MoMa Hub Modules
@@
In this section, we study the scaling behavior of MoMa to understand whether it benefits from a larger MoMa Hub. We first do a hub-scale ablation to progressively expand MoMa hub from 5 to 10 and 18 modules. Then we further expand MoMa Hub to include 12 QM9 modules (Ramakrishnan et al., 2014), which are trained on 12 quantum chemical properties for 134,000 stable small organic molecules. We consistently perform AMC and evaluation across all Hub variants. The full setup is described in Section C.6.
@@ Results
As presented in Table 3, as the MoMa hub scales, the average normalized test MAE across 17 tasks decreases monotonically (from 0.2040 with 5 modules to 0.1759 with 30 modules), showing no sign of saturation in this regime. The complete results are provided in Table 11 (Section D.5).
Table 3: Scaling with hub size. Average normalized test MAE decreases as the number of modules in MoMa Hub increases.
# Modules 5 10 18 30
Norm. MAE 0.2040 0.1910 0.1853 0.1759
To further analyze the effect of adding the 12 QM9 modules, we plot the test-MAE reduction rate against the AMC proxy-error decrease in Fig. 8 for datasets where QM9 modules are selected. We observe that: (1) The integration of QM9 modules leads to an average of 1.7% decrease in test set MAE; (2) a larger reduction in the AMC-optimized proxy error correlates with greater performance improvements post-fine-tuning (Pearson correlation =0.69=0.69). We highlight the task of MP Phonons prediction, which marks a 11.8% decrease in test set MAE after the inclusion of QM9 modules. Overall, these results support our vision of MoMa as a flexible community-driven platform: as more modules are added, downstream performance improves and AMC remains effective at larger scale.
Figure 7: Scatter plot showing the relationship between the test MAE decrease and the proxy error (Eq. 3) decrease after adding QM9 modules. The solid line represents a linear regression fit, yielding a Pearson correlation of 0.69.
Figure 8: Heat map of AMC weights on one data split. The x-axis represents the task names of the MoMa Hub modules, while the y-axis shows the 17 material tasks in Table 1. Darker color indicates a larger weight.
@@ 4.7 Materials Insights Mining
We argue that the AMC weights derived in Eq. 3 can provide valuable insights into the relationships of material properties. To explore this, we interpret the weights as indicators for the relationships between MoMa Hub modules and downstream tasks. Following Chang et al. (2022), we present a log-normalized visualization of these weights in Fig. 8.
@@ Results
We highlight several noteworthy observations: (1) The weights assigned by AMC effectively capture physically intuitive relationships between material properties. For instance, in predicting electronic dielectric constants, MoMa assigns high weights to the band gap modules, which is reasonable given the inverse relationship between the dielectric constant and the square of the band gap; (2) Less intuitive relationships also emerge. For the task of experimental band gap prediction (row 1), the DFT-computed formation energy module (column 1) is assigned the second-highest weight. In the prediction of dielectric constant (row 9), modules related to thermoelectric and thermal properties (columns 5 and 6) are heavily weighted. However, the first-principles relationship between these tasks is indirect. We hypothesize that in addition to task relevance, other factors such as data distribution and size may also influence the weight assignments for AMC.
5 Conclusion